References.

Abraham, W., 1979, ”Auβer”. In: H. Weydt, ed., Die Partikeln der deutschen

Sprache. De Gruyter, Berlin, pp. 239-255. Barwise, J. and R. Cooper, 1981, ”Generalized Quantifiers and Natural Language”.

Linguistics and Philosophy 4-1, pp. 159-219. Benthem, J. van, 1986, Essays in Logical Semantics. Reidel, Dordrecht. Chomsky, N., 1986, Barriers. MIT-Press, Cambridge, Massachusetts. Dowty, David, 1978, ”Lexically-Governed Transformations as Lexical Rules in a

Montague Grammar”, Linguistic Inquiry 9-3, pp. 393-426. Fiengo, R., J. Huang, H. Lasnik and T. Reinhart, 1988, ”The Syntax of Wh-in-Situ”, in:

H. Borer, ed., Proceedings of the Seventh West Coast Conference on Formal Linguistics, Stanford Linguistics Association, Stanford, pp. 81-98. Fintel, K. von, 1989, ”Exception Phrases”, talk presented at a conference on crosslinguistic quantification, Tucson.

Groenendijk, J. and M. Stokhof, 1984, Studies on the Semantics of Questions and the

Pragmatics of Answers. Doctoral dissertation, University of Amsterdam. Harris, Z., 1982, A Grammar of English on Mathematical Principles. John Wiley and

Sons, New York etc. Higginbotham, J. and R. May, 1981, ”Questions, Quantifiers and Crossing”, The

Linguistic Review 1, 41- 79. Hoeksema, J., 1983, ”Superlatieven”, Tabu 13, pp. 101-106. Hoeksema, J., 1986a, ”Monotonicity Phenomena in Natural Language”, Linguistic

Analysis 16, 1/2, pp. 25-40. Hoeksema, J., 1986b, ”Monotonie en Superlatieven”, in: C. Hoppenbrouwers, J.

Houtman, I. Schuurman and F. Zwarts, eds., Proeven van Taalwetenschap. Nederlands Instituut, Rijksuniversiteit Groningen, pp. 38-49.

Hoeksema, J., 1987, ”The Logic of Exception”, in: A. Miller and J. Powers, eds., Proceedings of the Fourth Eastern States Conference on Linguistics, The Ohio State University, Columbus, Ohio, pp. 100-113.

Hoeksema, J. and F. Zwarts, ”Some Remarks on Focus Adverbs”, to appear in the Journal of

Semantics. Horn, L. and S. Bayer, 1984, ”Short-Circuited Implicature: A Negative Contribution”,

Linguistics and Philosophy 7, pp. 397-414. Jackendoff, R. S., 1972, Semantic Interpretation in Generative Grammar, MIT-Press,

Cambridge, Massachusetts. Jacobs, J., 1983, Fokus und Skalen. Zur Syntax und Semantik der Grad-partikeln im

Deutschen. Niemeyer Verlag, Tübingen. Keenan, E.L. and Y. Stavi, 1986, ”A Semantic Characterization of Natural Language

Determiners”, Linguistics and Philosophy 9, pp. 253-326. König, E. and B. Kortmann, 1987, ”Absolute Complementation in the Lexical

Structure of English and German”, in: W.Lörscher and R. Schulze, eds., Perspectives on Language in Performance. Studies in Linguistics, Literary Criticism, and Language Teaching and Learning To Honour Werner Hüllen on the Occasion of His Sixtieth Birthday. Günther Narr, Tübingen, pp. 171-198.

Landman, F. and I. Moerdijk, 1980, ”Behalve als Voorzetsel”, Spektator 9-4, pp. 335-347.

Link, G., 1984, ”Hydras. On the Logic of Relative Constructions with Multiple Heads”, in: F. Landman and F. Veltman, eds., Varieties of Formal Semantics. Foris, Dordrecht, pp. 245-257.

Mayer, R., 1982, ”Korrektur, Spezifikation, Suspension und Widerspruch”, Linguistische

Berichte Papier Nr. 66 . McCarthy, J., 1980, ”Circumscription - A Form of Nonmonotonic Reasoning”, in:

Artificial Intelligence 13, pp. 27-39. Meijs, W., 1988, ”All but and if not in Brown and LOB”, in: M. Kytö, O. Ihalainen, and

M. Rissanen, eds., Corpus Linguistics, Hard and Soft. Proceedings of the Eighth International Conference on English Language Research on Computerized Corpora, Rodopi, Amsterdam, pp. 181-195.

Moignet, G., 1973, Les signes de l’exception dans l’histoire du Français. 2nd edition, Droz,

Geneva. Montague, R., 1973, ”The Proper Treatment of Quantification in Ordinary English”, in:

K.J.J. Hintikka, J. Moravcsik and P. Suppes, eds., Approaches to Natural Language. Reidel, Dordrecht.

Mourin, L., 1980, ”L’exception et la restriction dans les langues romanes”. Travaux de

Linguistique et de Litterature, vol. XVIII , 1, pp. 173-195. Neijt, A., 1979, Gapping. A Contribution to Sentence Grammar. Foris, Dordrecht. Paardekooper, P.C., 1966, ”Behalve als zn-patroondeel”, in: De Nieuwe Taalgids 59,

pp. 170-179. Ramsey, F.P., 1931, The Foundations of Mathematics and Other Logical Essays, Routledge

and Kegan Paul, London. Reinhart, T., 1989, ”Nonquantificational LF”, in A. Kasher, ed., The Chomskyan Turn. Russell, B., 1905, ”On Denoting”, Mind 14, pp. 479-493. Sag, I.A., G. Gazdar, T. Wasow and S. Weisler, 1985, ” Coordination and How to

Distinguish Categories”, Natural Language and Linguistic Theory 3-2, p. 117-171. Schmidt-Radefeldt, J., 1977, ”On So-Called ’Rhetorical Questions’”, in: Journal of

Pragmatics 1, pp. 375-392. Szabolcsi, A., 1985, ”Comparative Superlatives”, MIT Working Papers in Linguistics,

pp. 245-265. Vergnaud, J.-R., 1974, French Relative Clauses. Doctoral dissertation, MIT. Zwarts, F., 1983, ”Determiners: A Relational Perspective”, in: A.G.B. ter Meulen, ed.,

Studies in Modeltheoretic Semantics. Foris Publications, Dordrecht, pp. 37-62. Zwarts, F., 1985, ”De zaak vrijwel”. Tabu 15-4, pp. 145-149.